# Resource Management and Circuit Scheduling for Distributed Quantum Computing Interconnect Networks

### v3

@@ VII Conclusion
To achieve optimal assignment of QCircs to QPUs, we developed MILP formulations that jointly consider multiple objectives, including minimising inter-QPU infidelities and maximising QCirc throughput. These formulations also account for network topology, link latencies, QPU capacities, and QCirc structure, enabling effective and quantum-aware resource assignment. We evaluated both approaches through extensive simulations, showing that our methods support reliable QCirc execution while improving overall resource utilisation.
This work marks an initial step toward quantum-aware and network-aware job scheduling in cloud-based quantum data centres, incorporating key quantum-specific features such as qubit decoherence, as well as network-level considerations like topology. Several directions remain open for future research. First, our current model does not incorporate job priorities or deadlines. Future studies could develop more advanced scheduling strategies that address these aspects. Additionally, a detailed queue analysis under realistic quantum traffic models could provide deeper insights. Finally, leveraging machine learning techniques—particularly reinforcement learning—may enable the development of adaptive, scalable scheduling algorithms suited to complex quantum workloads.
This work provides a stepping stone toward scalable and intelligent resource management in future quantum cloud infrastructures.
@@ Acknowledgment
This work was supported by the Quantum Communication Hub funded by the EPSRC grant ref. EP/T001011/1 and the Quantum Data Centre of Future UKRI project (10004793). The authors would like to express their sincere gratitude to Prof. Reza Nejabati for his support and insightful technical discussions.

### v4

@@ V Conclusion
In this work, we address the challenge of efficient resource management and QCirc scheduling in distributed quantum computing networks. We propose two scheduling approaches: a dynamic Batch-QCirc scheduling method, which operates through sequential scheduling cycles with batch optimisation, and a Single-QCirc scheduling method that optimises QCirc allocation individually. To achieve optimal assignment of QCircs to QPUs, we developed MILP formulations that minimise errors arising from inter-QPU communication. These formulations also account for network topology, link loss, QPU capacities, and QCirc structure in the DQC, enabling effective and quantum resource-aware assignment. We combine both approaches with quantum circuit partitioning and evaluate their performance through extensive simulations, showing that our methods support reliable QCirc execution while improving overall resource utilisation.
This work takes an initial step toward quantum- and network-aware job scheduling in cloud-based quantum data centres, incorporating quantum-specific features and network considerations. Future work will explore priority- and time-aware scheduling, advanced learning models, and reinforcement learning for adaptive, scalable algorithms, paving the way for intelligent resource management in quantum cloud infrastructures.
@@ Acknowledgment
The authors would like to express their sincere gratitude to Reza Nejabati for his support and insightful technical discussions.